OpenStudy (anonymous):

what is half of 2pi/3?

OpenStudy (sanchez9457):

pi/3

OpenStudy (anonymous):

multiply it by 1/2

OpenStudy (unklerhaukus):

\[\frac12\times\frac{2\pi}3=\frac1{\cancel 2}\times\frac{\cancel2\pi}3\]